1. (6√2)/(√3)
2. √(18x^4y^3)

Answer by checkley77(12844)

1. (6√2)/(√3)
6*1.414/1.732
8.484/1.732=4.898 ans.
2.√(18x^4y^3)
sqrt(9*2x^2x^2y^2y) extract the sqrt9=3, the sqrtx^4=x^2 the sqrty^3=y
(3x^2y)sqrt(2y) ans.